OpenStudy (hang254):

Which equation is correct? sec x° = opposite ÷ adjacent cosec x° = opposite ÷ adjacent cosec x° = hypotenuse ÷ opposite sec x° = adjacent ÷ hypotenuse

OpenStudy (anonymous):

the thrid is correct..

OpenStudy (anonymous):

Here are the six ratios: \[\sin = \frac{opp.}{hyp.}\]\[\cos = \frac{adj.}{hyp.}\]\[\tan = \frac{opp.}{adj.}\]\[\csc = \frac{hyp.}{opp.}\]\[\sec = \frac{hyp.}{adj.}\]\[\cot = \frac{adj.}{opp.}\]

OpenStudy (hang254):

thank you

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):